OpenStudy (anonymous):

martha rakes leaves in 6 hours her brother in 7 hours how long will it take them together to do the job

OpenStudy (anonymous):

i came up with 42 im unsure

OpenStudy (dumbcow):

martha can do 1/6 of the job in 1 hour brother 1/7 of job in 1 hour 1/6 + 1/7 = 7/42 + 6/42 = 13/42 together they can do 13/42 of job in 1 hour 42/13 = 3 and 3/13 so it takes about 3 hours and 14 min

OpenStudy (anonymous):

someone help me

OpenStudy (anonymous):

i meant 42/13

OpenStudy (dumbcow):

you were correct :)

OpenStudy (anonymous):

thank you smartcow

OpenStudy (anonymous):

help u with what prettychick?